Calculate the atomic packing fraction of the fcc crystal structure.
1
Expert's answer
2019-02-03T08:31:07-0500

Calculate the atomic packing factor for FCC structure by the relation as follows:

no of atoms in a FCC structure =4

volume of four atoms = 4*(4πr^3)/3

volume of unit cell (cubic)=a^3

where r=a/(2√2)

Filling Factor = (4*(4πr^3)/3)/a^3 =((16πr^3)/3)/〖(2√2 r)〗^3 =√2/6 π=0.74